A car travels with an average velocity of 60km/h how long does it take to cover a distance of 500km
1
Expert's answer
2019-10-25T11:03:11-0400

The time will be


"t=\frac{s}{v}=\frac{500}{60}=8\frac{1}{3}\ h"

"t=8\ h\ 20\ min"